Jorge Company bottles and distributes B-Lite, a diet soft drink. The beverage is sold for 50 cents per 16-ounce bottle to retailers, who charge customers 78 cents per bottle. For the year 2014, management estimates the following revenues and costs.
Sales $1,804,000 Selling expenses—variable $69,800
Direct materials 428,000 Selling expenses—fixed 65,800
Direct labor 354,000 Administrative expenses—variable 64,920
Manufacturing overhead—variable 310,000 Administrative expenses—fixed 64,900
Manufacturing overhead—fixed 288,000
Jorge Company bottles and distributes B-Lite, a di
Jorge Company bottles and distributes B-Lite, a di
Prepare a CVP income statement for 2014 based on management’s estimates.

$

Jorge Company bottles and distributes B-Lite, a di
Jorge Company bottles and distributes B-Lite, a di
Calculate variable cost per bottle. (Round variable cost per bottle to 2 decimal places, e.g. 0.25.)
Variable cost per bottle

$

Jorge Company bottles and distributes B-Lite, a di
Jorge Company bottles and distributes B-Lite, a di
Compute the break-even point in (1) units and (2) dollars. (Round answers to 0 decimal places, e.g. 1,225.)
(1) Compute the break-even point

units
(2) Compute the break-even point

$

Jorge Company bottles and distributes B-Lite, a di
Jorge Company bottles and distributes B-Lite, a di
Compute the contribution margin ratio and the margin of safety ratio. (Round variable cost per bottle to 2 decimal places, e.g. 0.25 and final answers to 0 decimal places, e.g. 25%.)
Contribution margin ratio

%
Margin of safety ratio

%
Jorge Company bottles and distributes B-Lite, a di
Jorge Company bottles and distributes B-Lite, a di
Determine the sales dollars required to earn net income of $240,852. (Round answer to 0 decimal places, e.g. 1,225.)
Required sales dollars

$
Answer #1
Concepts and reason

CVP Analysis: Cost volume profit analysis is generally termed as CVP analysis. It determines the effect of operating income and net income due to the change in the cost and output. In CVP, variable cost per unit is constant, and total fixed cost is constant.

Breakeven analysis: This is one of the techniques of decision accounting which determines the particular level of activity where the total costs and the total revenues of a firm are equal. The relationship between cost, profit and sales volume provides the basis for the manager to take effective steps about the future profits. The break-even analysis uses the contribution margin approach.

Variable costs: Variable costs are direct costs that vary according to the business activities. It is generally referred to as unit-level costs which vary according to the number of units.

Direct material cost: Direct material cost is the cost related to the purchase of the raw materials that are directly related to the production of the goods. It includes opening stock of materials, purchases, cost of purchases, and deductions related to the closing stock of materials.

Direct labor cost: It refers to the cost of providing wages to the workers who are directly associated with the production of goods or services rendered to the customers. The cost of direct labor includes the wages, payroll taxes, and all the benefits sponsored by the manufacturer.

Overhead cost: Overhead costs are the indirect costs incurred to operate the business. They are not connected with the production of goods or rendering any service. They do not include any direct cost.

Fundamentals

Contribution margin income statement: Contribution margin income statement highlights the cost behaviour as variable costs and fixed costs and shows the contribution margin of the company.

Contribution margin: The difference between the selling price of revenue and the variable costs is called contribution margin. The contribution margin reports the revenue available to pay off fixed costs. The amount of revenue available after the fixed costs are paid off is the income or profit earned.

Contribution margin ratio: The contribution margin ratio is the ratio of contribution margin to total sales. It is calculated as given below:

Contributionmarginratio=ContributionmarginSales×100{\rm{Contribution margin ratio = }}\frac{{{\rm{Contribution margin}}}}{{{\rm{Sales}}}}{\rm{ }} \times {\rm{ 100}}

Break-even point: The break-even point is a point where revenue of the company equals the total cost incurred. In other words, the point of sales at which there is no profit or loss is called as Break-even Point (BEP). Hence, at the point of break-even, the profit of any company will be zero.

Sales revenue: Sales revenue is the total income earned by an organization by selling goods or rendering services.

Cost of goods sold: The costs that are incurred by a business to sell products in a particular period are called as cost of goods sold. It is also known as cost of sales. It is considered to be as the expense of the current period.

Fixed Cost: Fixed cost is a cost that remains same irrespective of the increase or decrease in the value of goods or any services rendered. It is the cost paid by the company that does not depend on the activities concerned with the business.

Contribution margin per unit: It is the excess of revenue per unit over the variable costs per unit. It is calculated as the difference between the sales per unit and variable costs per unit.

Margin of safety: Margin of safety is an excess of budgeted sales than break-even sales.

Margin of safety ratio: Margin of safety ratio is calculated by dividing the margin of safety by expected sales. It is also called as the margin of safety as a percentage of sales.

1)

Prepare a CVP income statement for 2014:

J company
CVP Income Statement
For the year ending December 31, 2014
Particulars Amount Amount
Sales
$ 1,804,000
Variable exp

Working note:

Calculate cost of goods sold:

Particulars
Direct material
Direct labor
Manufacturing overhead-variable
Cost of goods sold
Amount
$ 428,000
$ 354,000
$ 310,

2)

Calculate the variable cost per bottle:

Variablecostperbottle=TotalvariableexpensesNumberofunitssold=$1,226,720($1,804,00050%)=$0.34\begin{array}{c}\\{\rm{Variable cost per bottle = }}\frac{{{\rm{Total variable expenses}}}}{{{\rm{Number of units sold}}}}\\\\{\rm{ = }}\frac{{{\rm{\$ 1,226,720}}}}{{\left( {\frac{{{\rm{\$ 1,804,000}}}}{{{\rm{50\% }}}}} \right)}}\\\\{\rm{ = \$ 0}}{\rm{.34}}\\\end{array}

3)

Calculate the break-even point in units and the break-even point in dollars.

2,616,875 units
Compute the break-even point
(5418,700/($0.50-$0.34)
Compute the break-even point
(S2,616,875x$0.50)
$ 1,308,

4)

Calculate the contribution margin ratio:

Contributionmarginratio=ContributionmarginSales×100=$577,280$1,804,000×100=32%\begin{array}{c}\\{\rm{Contribution margin ratio = }}\frac{{{\rm{Contribution margin}}}}{{{\rm{Sales}}}}{\rm{ }} \times {\rm{ 100}}\\\\{\rm{ = }}\frac{{{\rm{\$ 577,280}}}}{{{\rm{\$ 1,804,000}}}}{\rm{ }} \times {\rm{ 100}}\\\\{\rm{ = 32\% }}\\\end{array}

Calculate the margin of safety ratio:

Marginofsafetyratio=((SalesBreakevensales)Sales)×100=(($1,804,000$1,308,438)$1,804,000)×100=27%\begin{array}{c}\\{\rm{Margin of safety ratio = }}\left( {\frac{{\left( {{\rm{Sales - Break even sales}}} \right)}}{{{\rm{Sales}}}}} \right) \times {\rm{100}}\\\\{\rm{ = }}\left( {\frac{{\left( {{\rm{\$ 1,804,000 - \$ 1,308,438}}} \right)}}{{{\rm{\$ 1,804,000}}}}} \right) \times {\rm{100}}\\\\{\rm{ = 27\% }}\\\end{array}

5)

Calculate the sales dollars required to earn net income of $240,852:

Requiredsalestoearnnetincomeof$240,852}=(Fixedcost+Targetednetincome)Contributionmarginratio=($418,700+$240,852)32%=$2,061,100\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Required sales to earn net}}\\\\{\rm{ income of \$ 240,852}}\\\end{array} \right\}{\rm{ = }}\frac{{\left( {{\rm{Fixed cost + Targeted net income}}} \right)}}{{{\rm{Contribution margin ratio}}}}\\\\{\rm{ = }}\frac{{\left( {{\rm{\$ 418,700 + \$ 240,852}}} \right)}}{{{\rm{32\% }}}}\\\\{\rm{ = \$ 2,061,100}}\\\end{array}

Ans: Part 1

J company
CVP Income Statement
For the year ending December 31, 2014
Particulars
Amount Amount
Sales
$ 1,804,000
Variable exp

Part 2

The variable cost per bottle is $0.34.

Part 3

| 1 Compute the break-even point
| 2 |Compute the break-even point
2,616,875 units
$ 1,308,438

Part 4

The contribution margin ratio is 32% and the margin of safety ratio is 27%.

Part 5

The required sales to earn net income of $240,852 are $2,061,100.